### v1

@@ 4.1 Infinite-Strategy Fictitious Play for Shared Effort Games
Given a shared effort game with players $N$, budget-defined strategies $S^{i}=\left\{x^{i}=(x^{i}_{\omega})_{\omega\in\Omega}\in\mathbb{R}_{+}^{\left|\Omega\right|}\mid\sum_{\omega\in\Omega_{i}}{x^{i}_{\omega}}\leq B_{i},\omega\notin\Omega_{i}\Rightarrow x^{i}_{\omega}=0\right\}$ and utilities $u^{i}(x)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\sum_{\omega\in\Omega_{i}}{\phi^{i}_{\omega}(x_{\omega})}$, define an Infinite-Strategy Fictitious Play (ISFP) as the following set of sequences. Consider a (pure) strategy in this game at time $1$, i.e. $(x^{i}(1))^{i\in N}=((x^{i}(1)_{\omega})_{\omega\in\Omega})^{i\in N}$, define $X^{i}(1)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left\{x^{i}(1)\right\}$, and define recursively, for each $i\in N$ and $t\geq 0$, the set of the possible strategic histories at time $t+1$ to be all the possible combinations of the history with the best responses to the previous strategic histories:
[…]
Here, the last play’s weight is $\frac{1}{\alpha t+1}$.
We do not know whether and when any convergence property can be proven for the generalised fictitious play in shared effort games. In simulations, our generalised fictitious play often converges to an NE, bolstering the importance of equilibrium efficiency, as stated by Roughgarden and Tardos (Nisan et al., 2007, Chapter 17).
4.2 Best Response
@@
In order to execute the infinite-strategy fictitious play, we from now on assume that the game has only two projects, $\Omega=\left\{\psi,\omega\right\}$. We would like to find a best response for a player $i\in N$, all the other players’ strategies $x^{-i}\in S^{-i}$ being fixed. From the weak monotonicity of the sharing functions $\phi^{i}_{\omega}$, we may assume w.l.o.g. that a best responding player contributes all her budget. Then, a strategy is uniquely determined by the contribution to project $\psi$ and we shall write $x^{i}$ for $x^{i}_{\psi}$, meaning that $x^{i}_{\omega}=B_{i}-x^{i}$.
Theorem 9.
4.3 The Simulation Method
[…]
4.4 Results and Conclusions
@@ 5 Mixed Nash Equilibrium
As we have seen, a game may not possess a pure NE. Therefore, we naturally turn to mixed extensions [^15] A mixed extension has strategies that are distributions on the pure strategies and the respective utilities are the expected utilities under these distributions. and ask whether a mixed extension always has an NE. At first, this is unclear. As the game is infinite, the theorem by Nash Nash (1951) about the existence of a mixed NE in finite games is irrelevant. Since the game is not continuous, even the theorem by Glicksberg Glicksberg (1952) about the existence of a mixed NE in continuous games is not applicable. Fortunately, we answer affirmatively employing a more general existence theorem by Maskin and Dasgupta (Maskin and Dasgupta, 1986, Theorem $5^{*}$), which requires some definitions. Let the strategy sets be $A_{i}\subset\mathbb{R}^{m}$. For each pair of players $i,j\in{1,\ldots,n}$, let $D(i)$ be a positive natural, and for a $d\in\left\{1,\ldots,D(i)\right\}$, let $f^{d}_{i,j}\colon\mathbb{R}\to\mathbb{R}$ be one-to-one and continuous, such that $(f^{d}_{i,j})^{-1}=f^{d}_{j,i}$. For every player $i$, we define
@@
Theorem $5^{*}$ from Maskin and Dasgupta (1986) states that if the strategy sets $A_{i}\subset\mathbb{R}^{m}$ are non-empty, convex, and compact, the utility function $u^{i}$ are continuous, except for $A^{**}(i)\subseteq A^{*}(i)$, the sum of all the utilities is upper semi-continuous, and for every player $i$, $u^{i}(a_{i},a_{-i})$ is bounded and weakly lower semi-continuous in $a_{i}$, then, there exists a mixed Nash equilibrium in this game.
Using Theorem $5^{*}$ from Maskin and Dasgupta (1986), we prove (see the proof in Section A)
@@
The existence result automatically extends to the solution concepts that include mixed Nash equilibria, such as correlated Aumann (1974) and coarse correlated Moulin and Vial (1978) equilibria. Luckily, not only existence results but also some efficiency bounds extend to other equilibria, as we describe next.
@@
As for Theorem 7 and its corollary, Roughgarden (Roughgarden, 2015, Theorem $3.2$) proves that the lower bounds on the price of anarchy that are proven using the smoothness argument go over to the mixed, correlated and coarse correlated equilibria.
@@ 6 Related Work
Motivation to contribute to public projects can be both extrinsic, like a payment or a record for the CV, and intrinsic, such as exercising one’s favorite skills Kaufmann et al. (2011) or, mostly, conservation citizen science projects Maund et al. (2020). Allahbakhsh et al. Allahbakhsh et al. (2013) discuss rewards for contributing to crowdsourcing, based on quality, where contributors have to fulfill given requirements and build reputation (extrinsic) over time. Wang et al. Wang and Fesenmaier (2003) model motivation to contribute to online travelling communities and conclude the importance of both the practical motives, such as supporting travellers, building relationships, and hoping for a future repay (extrinsic), as well as of internal drives to participate. Forte and Bruckman Forte and Bruckman (2005) study why people contribute to Wikipedia, by asking contributors, and conclude that the reasons are similar to those of scientists and include the desire to publish facts about the world (intrinsic). Bagnoli and Mckee Bagnoli and Mckee (1991) empirically check when people contribute to a public good, like building a playground. They find that if people know the threshold for the project’s success and benefit from collective contributing, then they will contribute, in agreement with the theory of Bagnoli and Lipman (1989). The work argues that knowing such information is realistic, suggesting the real cases of hiring a lobbyist and paying to a ski club as evidence. This conclusion supports our rationality assumption.
The concrete ways to motivate such contributions have been studied too. For instance, Harper et al. Harper et al. (2007) find that explicitly comparing a person’s contribution to the contribution of others helps focusing on the desired features of the system, but does not change the interest in the system per se. The influence of revealing how much people contribute to a movie rating community is experimentally studied in Chen et al. (2010). Initiating participation in online communities is experimentally studied in Ludford et al. (2004) on the example of the influence of similarity and uniqueness of ratings on participation.
@@
An NE can be inefficient, such as in the famous example of the prisoner’s dilemma (Osborne and Rubinstein, 1994, Example $16.2$). The ratios of the objectives in the least or the most efficient NE and in the optimum, called price of anarchy ($\operatorname{PoA}$Koutsoupias and Papadimitriou (1999) and price of stability ($\operatorname{PoS}$Anshelevich et al. (2004), respectively, constitute the most popular approaches to quantifying this inefficiency (Nisan et al., 2007, Chapter 17). The price of anarchy measures the best guarantee on an NE, while the price of stability measures the cost of leading the game to a specific equilibrium. Roughgarden and Tardos (Nisan et al., 2007, Chapter 17) discuss inefficiency of equilibria in non-cooperative games and consider the examples of network, load balancing and resource allocation games. This work argues that understanding exactly when selfish behaviour is socially profitable is important, since in many applications, implementing control is extremely difficult.
@@
In our simulations, we generalise and employ the widely studied fictitious play, introduced by Brown Brown (1951), to find equilibria. In this play, each player best-responds to the product of cumulative marginal histories of the others’ actions at every time step. It is a myopic learning process. If the game is finite, then if a fictitious play converges, then the distribution in its limit is an NE Mannor (2008). Conversely, a game is said to possess the fictitious play property if every fictitious play approaches equilibrium in this game Monderer and Shapley (1996a). Many researchers show games that possess this property, for example, finite two-player zero-sum games Robinson (1951) and finite weighted potential games Monderer and Shapley (1996a). A famous example for a game without such property is a $3\times 3$ game by Shapley Shapley et al. (1964). In this game, there is a cyclic fictitious play that plays each strategy profile for at least an exponentially growing number of times, and therefore, does not converge at all. Moreover, even its subsequences do not converge to an NE.
@@
Since pure equilibria may not exist, we also consider the mixed extensions of a game, where the strategies are the probability distributions over the original (pure) strategies and the utilities are defined as the expected utilities under the played distributions. Then, a mixed Nash equilibrium is defined as an equilibrium of the mixed extension of a game. Regarding the existence of a mixed NE, Nash proved in his classical paper Nash (1951) that a mixed extension of a finite game always possesses an NE. Glikcsberg Glicksberg (1952) showed the existence of a mixed NE for continuous games. However, in shared effort games with a positive threshold, the threshold creates discontinuity. Dasgupta and Maskin prove the existence of a mixed equilibrium for a subclass of possibly discontinuous games Maskin and Dasgupta (1986). We show that shared effort games can be cast to fit Dasgupta and Maskin in Theorem 10.
@@
The models that resemble ours, but expressed in the terms of cooperative games, where every coalition of players has a value, include the works by Zick, Elkind and Chalkiadakis Zick and Elkind (2011); Zick et al. (2012). Contributing to a coalition can be considered as contributing to a project in a shared effort game, and in both cases players have budgets and obtain shares. Another coalitional model, where participating in a coalition can be seen as contributing to projects appears in Gal et al. (2020), though in addition to being a different kind of game, this model allows no participation in multiple coalitions. Despite these similarities, cooperative games are not games in our sense, since they do not define utilities. Moreover, even if we considered a shared effort game as a particular case of a cooperative game, a positive threshold would vitiate the individual rationality of the core, since a player who obtains a positive share when she is the only contributor to a project may obtain nothing when others contribute to the same project.
@@
We now move to non-cooperative games modelling sharing effort. Shared effort games where only the biggest contributor obtains the project’s value, while everyone pays, are called all-pay auctions, and their equilibria are studied, for instance, by Baye, Kovenock and de Vries Michael R. Baye (1996). That work shows cases where each player obtains the expected payoff of zero, and cases where the winner obtains the difference between the two highest valuations, while the rest obtains zero. All-pay auctions model lobbying, single-winner contests, political campaigns, striving for a job promotion (see e.g. Siegel (2009)) and Colonel Blotto games with two Roberson (2006) or more Kovenock and Roberson (2021); Boix-Adserà et al. (2021) players. In a Colonel Blotto game, the colonels divide their armies between battlefields, and at every battlefield, the larger force wins. The number of the won battlefields defines the utility of a colonel. Namely, the winner of a battlefield takes it fully to her utility. For two players, Roberson Roberson (2006) analyses the equilibria of this game and their expected payoffs. Any outcome is socially optimum, since this is a constant-sum game, modelling a confrontation. The $n$-player Blotto game has been studied too Kovenock and Roberson (2021). Shared effort games can model these games, but typically, shared effort models cooperation, where a project’s value increases with the total contribution.
@@
The models below feature no thresholds required to share a project’s value. Anshelevich and Hoefer (Anshelevich and Hoefer, 2012) consider an undirected graph model, where the nodes are the players and each player divides its budget between its adjacent edges in minimum effort games (where the edges are $2$-player projects), each of which equally rewards both sides by the value of the project’s success (i.e., duplication instead of division). Anshelevich and Hoefer prove the existence of equilibria, find the complexity of finding an NE, and find that the $\operatorname{PoA}$ is at most $2$. A related setting of multi-party computation games appears in Smorodinsky and Tennenholtz Smorodinsky and Tennenholtz (2006). There, the players are computing a common function that requires them to compute a costly private value, motivating free-riding. The work suggests a mechanism, where honest computation is an NE. This differs from our work, since they consider cost minimization, and the choice of the players is either honestly computing or free riding, without choosing projects. For shared effort games with specific conditions (obtaining at least a constant share of one’s marginal contribution to the project’s value and no contribution threshold, i.e. $\theta=0$), Bachrach et al. (Bachrach et al., 2011) show that the price of anarchy ($\operatorname{PoA}$) is at most the number of players. That work also upper bounds the $\operatorname{PoA}$ for the case of convex project functions, where each player receives at least a constant share of its marginal contribution to the project’s value. The academic game Ackerman and Branzei (2017) considers discrete weights, limits the collaboration to at most two agents per project, and the utility functions model synergy and various ordering of the authors, which effect is studied. This model employs the pairwise stability approach Jackson and Wolinsky (1996), which in addition to guarding against unilateral deviations, guards against $2$ agents deviating to a common project too. Gollapudi et al. Gollapudi et al. (2017) allow an agent merely to choose which single team to join. They consider several profit sharing models, including equal sharing, but without any threshold.
@@
There has been no research of the NE of our problem with a $\theta\in[0,1]$ sharing mechanism before the preliminary version Polevoy et al. (2014) of our paper. A positive $\theta\in(0,1]$ contradicts the condition of receiving a constant share. We provide precise conditions for existence of NE, and find their efficiency. There exist two works that variate this model in various directions. In Polevoy and de Weerdt (2018a), quotas and other requirements from projects, such as paper and grants applications, are modeled by requiring a project to be good enough in order to actually obtain its value. They find equilibria and their efficiency and compare the efficiency of equilibria for being within a quota and for attaining at least a certain minimum value. In Polevoy and de Weerdt (2018b), investing in reciprocal interactions, like attending seminars and meeting friends, is modeled be assuming each project to be a reciprocal interaction from which the interacting parties obtain a value. They prove the convergence of such a process and analyse the existence and efficiency of the equilibria.

### v2

@@ 4 Narrowing down Nash to Cyclically Strong Equilibrium
A cyclically strong equilibrium is a Nash equilibrium $x$, where there exists no coalitional deviation of players $i_{1},i_{2},\ldots,i_{p}$ such that
  1. 1.
    for each $l=1,2,\ldots,p$, there exists a project $\omega\in\Omega$, such that $x^{i_{l}}_{\omega}=B_{i_{l}}$, namely, all $i_{l}$’s budget is in one project;
  2. 2.
    if each player $i_{l}$ above deviates to invest all $B_{i_{l}}$ into the cyclically next player’s project, thereby creating profile $x^{\prime}$, then no deviating player loses and at least one of them strictly benefits w.r.t. $x$.
Cyclically strong equilibrium is a restriction of Nash and a relaxation of strong Nash equilibrium [6].
Interestingly, the new concept often coincides with Nash equilibrium.
@@
In any thresholded shared effort game where $B_{i}\geq\theta B_{j},\forall i,j\in N$, any Nash equilibrium is also cyclically strong.
Proof. Consider an $x\in\text{NE}$, and consider a cyclic coalitional deviation by players $i_{1},i_{2},\ldots,i_{p}$ from Definition 4, assuming by contradiction that no player loses and at least one strictly benefits. Since no unilateral deviation is profitable, while now no player loses and at least one strictly benefits, then either some $B_{i_{l}}<\theta B_{i_{l+1}}$ (modulo $p$), which is assumed not to be the case, or each deviation on this cycle has to be of a budget of size at most the size of the previous deviation, thus $B_{i_{1}}\geq B_{i_{2}}\geq\ldots\geq B_{i_{p}}\geq B_{i_{1}}$, with at least one inequality being strict. However, this cycle implies $B_{i_{1}}>B_{i_{1}}$, in contradiction to the assumption of a cyclic deviation above. ∎
This property allows for a better theoretical study of Nash equilibria, as well as providing a practical property they all possess.
@@ 5 Mixed Nash Equilibrium
As we have seen, a game may not possess a pure NE. Therefore, we naturally turn to mixed extensions [^13] A mixed extension has strategies that are distributions on the pure strategies and the respective utilities are the expected utilities under these distributions. and ask whether a mixed extension always has an NE. At first, this is unclear. As the game is infinite, the theorem by Nash [38] about the existence of a mixed NE in finite games is irrelevant. Since the game is not continuous, even the theorem by Glicksberg [18] about the existence of a mixed NE in continuous games is not applicable. Fortunately, we answer affirmatively employing a more general existence theorem by Maskin and Dasgupta [32, Theorem $5^{*}$], which requires some definitions. Let the strategy sets be $A_{i}\subset\mathbb{R}^{m}$. For each pair of players $i,j\in{1,\ldots,n}$, let $D(i)$ be a positive natural, and for a $d\in\left\{1,\ldots,D(i)\right\}$, let $f^{d}_{i,j}\colon\mathbb{R}\to\mathbb{R}$ be one-to-one and continuous, such that $(f^{d}_{i,j})^{-1}=f^{d}_{j,i}$. For every player $i$, we define
@@
Theorem $5^{*}$ from [32] states that if the strategy sets $A_{i}\subset\mathbb{R}^{m}$ are non-empty, convex, and compact, the utility function $u^{i}$ are continuous, except for $A^{**}(i)\subseteq A^{*}(i)$, the sum of all the utilities is upper semi-continuous, and for every player $i$, $u^{i}(a_{i},a_{-i})$ is bounded and weakly lower semi-continuous in $a_{i}$, then, there exists a mixed Nash equilibrium in this game.
Using Theorem $5^{*}$ from [32], we prove (see the proof in Section A)
@@
The existence result automatically extends to the solution concepts that include mixed Nash equilibria, such as correlated [7] and coarse correlated [37] equilibria. Luckily, not only existence results but also some efficiency bounds extend to other equilibria, as we describe next.
@@
As for Theorem 7 and its corollary, Roughgarden [46, Theorem $3.2$] proves that the lower bounds on the price of anarchy that are proven using the smoothness argument go over to the mixed, correlated and coarse correlated equilibria.
@@ 6 Related Work
Motivation to contribute to public projects can be both extrinsic, like a payment or a record for the CV, and intrinsic, such as exercising one’s favorite skills [23] or, mostly, conservation citizen science projects [33]. Allahbakhsh et al. [3] discuss rewards for contributing to crowdsourcing, based on quality, where contributors have to fulfill given requirements and build reputation (extrinsic) over time. Wang et al. [51] model motivation to contribute to online travelling communities and conclude the importance of both the practical motives, such as supporting travellers, building relationships, and hoping for a future repay (extrinsic), as well as of internal drives to participate. Forte and Bruckman [16] study why people contribute to Wikipedia, by asking contributors, and conclude that the reasons are similar to those of scientists and include the desire to publish facts about the world (intrinsic). Bagnoli and Mckee [10] empirically check when people contribute to a public good, like building a playground. They find that if people know the threshold for the project’s success and benefit from collective contributing, then they will contribute, in agreement with the theory of [9]. The work argues that knowing such information is realistic, suggesting the real cases of hiring a lobbyist and paying to a ski club as evidence. This conclusion supports our rationality assumption.
The concrete ways to motivate such contributions have been studied too. For instance, Harper et al. [21] find that explicitly comparing a person’s contribution to the contribution of others helps focusing on the desired features of the system, but does not change the interest in the system per se. The influence of revealing how much people contribute to a movie rating community is experimentally studied in [13]. Initiating participation in online communities is experimentally studied in [29] on the example of the influence of similarity and uniqueness of ratings on participation.
@@
An NE can be inefficient, such as in the famous example of the prisoner’s dilemma [40, Example $16.2$]. The ratios of the objectives in the least or the most efficient NE and in the optimum, called price of anarchy ($\operatorname{PoA}$[26] and price of stability ($\operatorname{PoS}$[4], respectively, constitute the most popular approaches to quantifying this inefficiency [39, Chapter 17]. The price of anarchy measures the best guarantee on an NE, while the price of stability measures the cost of leading the game to a specific equilibrium. Roughgarden and Tardos [39, Chapter 17] discuss inefficiency of equilibria in non-cooperative games and consider the examples of network, load balancing and resource allocation games. This work argues that understanding exactly when selfish behaviour is socially profitable is important, since in many applications, implementing control is extremely difficult.
@@
In our simulations, we generalise and employ the widely studied fictitious play, introduced by Brown [12], to find equilibria. In this play, each player best-responds to the product of cumulative marginal histories of the others’ actions at every time step. It is a myopic learning process. If the game is finite, then if a fictitious play converges, then the distribution in its limit is an NE [30]. Conversely, a game is said to possess the fictitious play property if every fictitious play approaches equilibrium in this game [35]. Many researchers show games that possess this property, for example, finite two-player zero-sum games [45] and finite weighted potential games [35]. A famous example for a game without such property is a $3\times 3$ game by Shapley [48]. In this game, there is a cyclic fictitious play that plays each strategy profile for at least an exponentially growing number of times, and therefore, does not converge at all. Moreover, even its subsequences do not converge to an NE.
@@
Since pure equilibria may not exist, we also consider the mixed extensions of a game, where the strategies are the probability distributions over the original (pure) strategies and the utilities are defined as the expected utilities under the played distributions. Then, a mixed Nash equilibrium is defined as an equilibrium of the mixed extension of a game. Regarding the existence of a mixed NE, Nash proved in his classical paper [38] that a mixed extension of a finite game always possesses an NE. Glikcsberg [18] showed the existence of a mixed NE for continuous games. However, in shared effort games with a positive threshold, the threshold creates discontinuity. Dasgupta and Maskin prove the existence of a mixed equilibrium for a subclass of possibly discontinuous games [32]. We show that shared effort games can be cast to fit Dasgupta and Maskin in Theorem 9.
@@
The models that resemble ours, but expressed in the terms of cooperative games, where every coalition of players has a value, include the works by Zick, Elkind and Chalkiadakis [53, 52]. Contributing to a coalition can be considered as contributing to a project in a shared effort game, and in both cases players have budgets and obtain shares. Another coalitional model, where participating in a coalition can be seen as contributing to projects appears in [17], though in addition to being a different kind of game, this model allows no participation in multiple coalitions. Despite these similarities, cooperative games are not games in our sense, since they do not define utilities. Moreover, even if we considered a shared effort game as a particular case of a cooperative game, a positive threshold would vitiate the individual rationality of the core, since a player who obtains a positive share when she is the only contributor to a project may obtain nothing when others contribute to the same project.
@@
We now move to non-cooperative games modelling sharing effort. Shared effort games where only the biggest contributor obtains the project’s value, while everyone pays, are called all-pay auctions, and their equilibria are studied, for instance, by Baye, Kovenock and de Vries [34]. That work shows cases where each player obtains the expected payoff of zero, and cases where the winner obtains the difference between the two highest valuations, while the rest obtains zero. All-pay auctions model lobbying, single-winner contests, political campaigns, striving for a job promotion (see e.g. [49]) and Colonel Blotto games with two [44] or more [28, 11] players. In a Colonel Blotto game, the colonels divide their armies between battlefields, and at every battlefield, the larger force wins. The number of the won battlefields defines the utility of a colonel. Namely, the winner of a battlefield takes it fully to her utility. For two players, Roberson [44] analyses the equilibria of this game and their expected payoffs. Any outcome is socially optimum, since this is a constant-sum game, modelling a confrontation. The $n$-player Blotto game has been studied too [28]. Shared effort games can model these games, but typically, shared effort models cooperation, where a project’s value increases with the total contribution.
@@
The models below feature no thresholds required to share a project’s value. Anshelevich and Hoefer [5] consider an undirected graph model, where the nodes are the players and each player divides its budget between its adjacent edges in minimum effort games (where the edges are $2$-player projects), each of which equally rewards both sides by the value of the project’s success (i.e., duplication instead of division). Anshelevich and Hoefer prove the existence of equilibria, find the complexity of finding an NE, and find that the $\operatorname{PoA}$ is at most $2$. A related setting of multi-party computation games appears in Smorodinsky and Tennenholtz [50]. There, the players are computing a common function that requires them to compute a costly private value, motivating free-riding. The work suggests a mechanism, where honest computation is an NE. This differs from our work, since they consider cost minimization, and the choice of the players is either honestly computing or free riding, without choosing projects. For shared effort games with specific conditions (obtaining at least a constant share of one’s marginal contribution to the project’s value and no contribution threshold, i.e. $\theta=0$), Bachrach et al. [8] show that the price of anarchy ($\operatorname{PoA}$) is at most the number of players. That work also upper bounds the $\operatorname{PoA}$ for the case of convex project functions, where each player receives at least a constant share of its marginal contribution to the project’s value. The academic game [2] considers discrete weights, limits the collaboration to at most two agents per project, and the utility functions model synergy and various ordering of the authors, which effect is studied. This model employs the pairwise stability approach [22], which in addition to guarding against unilateral deviations, guards against $2$ agents deviating to a common project too. Gollapudi et al. [19] allow an agent merely to choose which single team to join. They consider several profit sharing models, including equal sharing, but without any threshold.
@@
There has been no research of the NE of our problem with a $\theta\in[0,1]$ sharing mechanism before the preliminary version [43] of our paper. A positive $\theta\in(0,1]$ contradicts the condition of receiving a constant share. We provide precise conditions for existence of NE, and find their efficiency. There exist two works that variate this model in various directions. In [42], quotas and other requirements from projects, such as paper and grants applications, are modeled by requiring a project to be good enough in order to actually obtain its value. They find equilibria and their efficiency and compare the efficiency of equilibria for being within a quota and for attaining at least a certain minimum value. In [41], investing in reciprocal interactions, like attending seminars and meeting friends, is modeled be assuming each project to be a reciprocal interaction from which the interacting parties obtain a value. They prove the convergence of such a process and analyse the existence and efficiency of the equilibria.